Manufacturing cost data for Orlando Company, which uses a job order cost system, are presented below.

Indicate the missing amount for each letter. Assume that in all cases manufacturing overhead is applied on the basis of direct labor cost and the rate is the same. (Round overhead rate to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 5,275.)

Case A

Case B

Case C

Direct materials used $

Entry field with correct answer

(a) $86,200 $68,150
Direct labor 52,000 144,800

Entry field with incorrect answer

(h)
Manufacturing overhead applied 39,000

Entry field with incorrect answer

(d)

Entry field with incorrect answer

(i)
Total manufacturing costs 156,550

Entry field with incorrect answer

(e) 216,900
Work in process 1/1/17

Entry field with correct answer

(b) 23,100 18,400
Total cost of work in process 203,700

Entry field with incorrect answer

(f)

Entry field with incorrect answer

(j)
Work in process 12/31/17

Entry field with incorrect answer

(c) 15,400

Entry field with incorrect answer

(k)
Cost of goods manufactured 199,000

Entry field with incorrect answer

(g) 224,800
Answer #1

Answer:

Case A Case B Case C
Amount($) Amount($) Amount($)
A Direct materials used                   65,550 (a)        86,200           68,150
B Direct labor                   52,000     1,44,800           85,000 h
C Manufacturing overhead applied                   39,000     1,08,600 d           63,750 i
D=A+B+C Total manufacturing costs                1,56,550     3,39,600 e        2,16,900
E Work in process 1/1/17                   47,150 b        23,100           18,400
F=D+E Total cost of work in process                2,03,700     3,62,700 f        2,35,300 j
G Work in process 12/31/17                     4,700 c        15,400           10,500 k
I=F-G Cost of goods manufactured                1,99,000     3,47,300 g        2,24,800
Amount($)
a               65,550 156550-39000-52000
b               47,150 203700-156550
c                 4,700 203700-199000
d            1,08,600 144800*(39000/52000)
e            3,39,600 86200+144800+108600
f            3,62,700 339600+23100
g            3,47,300 362700-15400
h               85,000 (216900-68150)/(1+39000/52000)
i               63,750 85000*(39000/52000)
j            2,35,300 216900+18400
k               10,500 235600-224800